# Subdiffusive concentration for the chemical distance in Bernoulli percolation

### v1

@@ Abstract.
Considering supercritical Bernoulli percolation on ${\mathbb{Z}}^{d}$, Garet and Marchand [GM09] proved a diffusive concentration for the graph distance. In this paper, we sharpen this result by establishing the subdiffusive concentration inequality, which revisits the sublinear bound of the variance proved by Dembin [Dem22] as a consequence. Our approach is inspired by similar work in First-passage percolation [BR08, DHS14], combined with new tools to address the challenge posed by the infinite weight of the model. These tools, including the notion of effective radius and its properties, enable a simple one-step renormalization process as a systematic means of managing the effects of resampling edges.
V.H. Can: Institute of Mathematics, Vietnam Academy of Science and Technology, 18 Hoang Quoc Viet, Cau Giay, Hanoi, Vietnam. cvhao@math.ac.vn
V.Q. Nguyen: Institute of Mathematics, Vietnam Academy of Science and Technology, 18 Hoang Quoc Viet, Cau Giay, Hanoi, Vietnam. nvquyet@math.ac.vn
@@ 1.1. Model and main result
Bernoulli percolation is a simple but well-known probabilistic model for porous material introduced by Broadbent and Hammersley [BH57]. Let $d\geq 2$ and ${{\mathcal{E}}({\mathbb{Z}}^{d})}$ be the set of the edges $e=(x,y)$ of endpoints $x=(x_{1},\ldots,x_{d}),y=(y_{1},\ldots,y_{d})\in{\mathbb{Z}}^{d}$ such that $\|x-y\|_{1}:=\sum_{i=1}^{d}|x_{i}-y_{i}|=1$. Given the parameter $p\in(0,1)$, we let each edge $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ be open with probability $p$ and closed otherwise, independently of the state of other edges. The phase transition of the model has been well-known since 1960s. In particular, it has been shown that there exists a critical parameter $p_{c}(d)\in(0,1)$ such that there is almost surely a unique infinite open cluster ${\mathcal{C}}_{\infty}$ if $p>p_{c}(d)$, whereas all open clusters are finite if $p<p_{c}(d)$. We refer to the book of Grimmett [Gri89] and the reviews of Duminil-Copin [DC18] for classical results and recent developments in percolation theory. In this paper, we are interested in the chemical distance that is the graph distance of infinite cluster in the supercritical regime.
@@
${\mathrm{D}}^{U}(A,B)$ $:=\inf\{|\gamma|:x\in A,\,y\in B,\,\gamma\text{ is an open nearest-neighbor path from }x\text{ to }y\text{ inside }U\}.$
When $U={\mathbb{Z}}^{d}$, we simply write ${\mathrm{D}}$ for ${\mathrm{D}}^{{\mathbb{Z}}^{d}}$ and we write ${\mathrm{D}}(x,y)$ for ${\mathrm{D}}(\{x\},\{y\})$. Since this graph distance between two disconnected vertices is not well defined, we consider a regularized version as follows. Let $x\in{\mathbb{Z}}^{d}$, we denote by $x^{*}$ the closest point to $x$ in ${\mathcal{C}}_{\infty}$ (in $\|.\|_{\infty}$ distance) and call it the regularized point of $x$. The graph distance is then defined as: for $x,y\in{\mathbb{Z}}^{d}$,
${\mathrm{D}}^{*}(x,y):={\mathrm{D}}(x^{*},y^{*})=\inf\{|\gamma|:\gamma\text{ is an open nearest-neighbor path from }x^{*}\text{ to }y^{*}\}.$
@@
The linear growth of ${\mathrm{D}}^{*}_{n}$ was described by Garet and Marchand [GM04]: for any $p>p_{c}(d)$, there exists a constant $\mu_{p}({\mathbf{e}}_{1})\in(0,\infty)$ such that,
$\lim_{n\to\infty}\dfrac{{\mathrm{D}}^{*}_{n}}{n}=\mu_{p}({\mathbf{e}}_{1})\quad\text{a.e and in }L^{1}.$ (1.1)
The value $\mu_{p}({\mathbf{e}}_{1})$ is called the time constant. The regularity of the time constant (Lipschitz-continuity of the time constant with respect to $p$) was proved in [CD22] and revisited in [CNN24]. Naturally, the next question we are interested in is the fluctuation and deviation of the graph distance. The large deviation of ${\mathrm{D}}^{*}_{n}$ has been well studied since 1990s, see, for example, [AP96, GM07, DN22]. The moderate deviation of ${\mathrm{D}}^{*}_{n}$ (or precisely the concentration with diffusive scale) was established by Garet and Marchand (see [GM09, Theorem 1.2]), using concentration inequalities of Boucheron, Lugosi, and Massart: for each $C>0$, there exists some constant $c>0$ such that for all $\lambda\in[C\log n,\sqrt{n}$],
$\mathbb{P}[|{\mathrm{D}}^{*}_{n}-{\mathbb{E}}[{\mathrm{D}}^{*}_{n}]|\geq\sqrt{n}\lambda]\leq c^{-1}e^{-c\lambda}.$ (1.2)
Recently, Dembin gave a sublinear bound on the vairance of ${\mathrm{D}}^{*}_{n}$ (see [Dem22, Theorem 1.1]): there exists a positive constant $C$ such that
$\operatorname{Var}({\mathrm{D}}^{*}_{n})\leq\frac{Cn}{\log n}.$ (1.3)
The main result of our paper is to sharpen the moderate deviation (1.2) by establishing a sub-diffusive concentration of ${\mathrm{D}}^{*}_{n}$.
@@
Let $p>p_{c}(d)$. There exists a positive constant $c>0$ such that for all $n\in{\mathbb{N}}$ and $\kappa\geq 0$,
$\mathbb{P}\left(|{\mathrm{D}}^{*}_{n}-\mathbb{E}[{\mathrm{D}}^{*}_{n}]|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}e^{-c\kappa}.$ (1.4)
Consequently, the sublinear bound of the variance (1.3) holds.
@@
If $\kappa<1$, we can take $c$ small enough such that (1.4) holds trivially. From now on, we focus on the case $\kappa\geq 1$ throughout this paper.
@@
${\mathrm{T}}^{U}(x,y):=\inf_{\begin{subarray}{c}\gamma:x\to y\\ \gamma\subset U\end{subarray}}\sum_{e\in\gamma}t_{e},$ (1.5)
@@
${\mathrm{T}}^{*}(x,y):={\mathrm{T}}(x^{*},y^{*}),$ (1.6)
where $z^{*}$ denotes the closest point to $z$ in the infinite cluster of finite weight edges. Then the supercritical Bernoulli percolation can be referred as a particular case of FPP with distribution $\zeta=\zeta_{p}=p\delta_{1}+(1-p)\delta_{\infty}$.
The convergence of the scaled passage time in probability to time constant was obtained by Cerf and Théret [CT16, Theorem 4], without any moment assumption: there exists a constant $\mu_{\zeta}(\mathbf{e}_{1})\in[0,\infty)$ such that
$\lim_{n\to\infty}\frac{{\mathrm{T}}^{*}(0,n\mathbf{e}_{1})}{n}=\mu_{\zeta}(\mathbf{e}_{1})\qquad\text{ in probability}.$ (1.7)
In [GM04, Remark 1], Garet and Marchand proved that if ${\mathbb{E}}[t_{e}^{2+\varepsilon}{\mathbb{I}}(t_{e}<\infty)]<\infty$ with some $\varepsilon>0$, then the convergence in (1.7) holds true almost surely and in $L_{1}$. The continuity property of the time constant with respect to $\zeta$ was first shown by Cox and Kesten in [Cox80, CK81, Kes86] for FPP with finite weights. This was later extended for FPP with possibly infinite weights by Garet, Marchand, Procaccia, and Théret in [GMPT17]. It then has been proved by Damron, Hanson, and Sosoe in [DHS15] that if $\zeta([0,\infty))=1$ and ${\mathbb{E}}[t_{e}^{2}\log_{+}t_{e}]<\infty$, then the sublinear variance holds:
$\operatorname{Var}({\mathrm{T}}(0,n{\mathbf{e}}_{1}))\leq Cn/\log n.$ (1.8)
The sublinearity of variance is a particular case of superconcentration, a phenomenon coined by Chatterjee [Cha14] describing the situation that usual concentration tools such as Poicaré or Erfon-Stein inequality give suboptimal bounds. In fact, in some particular models as Gaussian disordered systems [Cha14], he also established a connection between superconcentration and a chaotic phenomenon of the ground states. However, understanding these notions beyond the Gaussian realm remains quite limited. Recently, Ahlberg, Deijfen, and Sfragara [ADS23] showed that this deep relation holds true in the context of First-passage percolation, under the assumption that $\zeta$ is finitely supported and has a finite second moment. Other examples of the superconcentration phenomena can be found in [BDG20, Cha23, DG24].
@@ 1.2. Method of the proof
The general strategy of proving the subdiffusive concentration for FPP has been shown in [BR08] and [DHS14]. First, by general bounds, the subdiffusive concentration of ${\mathrm{T}}(0,n{\mathbf{e}}_{1})$ can be reduced to some estimates of the variance of exponential functionals of ${\mathrm{T}}(0,n{\mathbf{e}}_{1})$. Next, to bound the variance of these exponential terms, one could use the entropy inequalities combining with the geometric averaging trick of Benjamini, Kalai, Schramm introduced in their highly influential paper [BKS03]. Moreover, in FPP [DHS14], the key for variance or entropy bounds is to well control for the effect of resampling an edge weight, and the moment conditions play a crucial role in these controls. However, in the context of the graph distance in Bernoulli percolation, closing an edge on the geodesic can have a significant impact on the graph distance due to the possibility of infinite edge-weight values. By introducing the notion effective radius and exploiting the greedy lattice animals, we provide a simple one-step renormalization process as a systematic way to manage the effect of resampling edges. Let us now explain our proof in more detail.
@@
There exists a positive constant $c$ such that for all $n\in{\mathbb{N}}$ and $\kappa\geq 0$
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-\mathbb{E}[{{\mathrm{T}}_{n}}]|\geq\sqrt{\tfrac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa).$ (1.9)
Our argument initially follows the common scheme as in [BR08, DHS14]. Considering ${\mathrm{F}}_{m}$ a spatial average of ${\mathrm{T}}_{n}$ defined in (4.3) (inspired by Benjamini, Kalai, and Schramm in [BKS03]), the proof of (1.9) can be transfered to the following. There exists a constant $c>0$ such that with $K=\dfrac{cn}{\log n}$,
$\forall\,|\lambda|<\dfrac{1}{\sqrt{K}},\quad\operatorname{Var}\left[e^{\lambda{{\mathrm{F}}_{m}}}\right]\leq K\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]<\infty.$ (1.10)
Using Falik-Samorodnitsky inequality (Lemma 4.1), the problem (1.10) is then reduced to understand well the behavior of $G=e^{\lambda{\mathrm{F}}_{m}}$ when flipping the state of edges in ${{\mathcal{E}}({\mathbb{Z}}^{d})}$. In fact, we shall show in Section 4 that the above variance bound is closely related to a large deviation estimate
${\mathbb{P}}\Bigg{(}\sum_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}(\nabla_{e}{\mathrm{T}}_{n})^{2}\geq Cn\Bigg{)}\leq\exp(-\sqrt{n}(\log n)^{C}),\qquad\nabla_{e}{\mathrm{T}}_{n}={\mathrm{T}}_{n}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{n}(1,t_{e^{c}}),$ (1.11)
@@
$|\nabla_{e}{\mathrm{T}}_{n}|\leq\hat{R}_{e}{\mathbb{I}}(e\in\gamma),\quad\hat{R}_{e}=\min\{C_{*}R_{e},\log^{2}n\},$ (1.12)
with $C_{*}$ being a constant introduced in (2.2) and $\gamma$ is a geodesic of ${\mathrm{T}}_{n}$. While the bound by $C_{*}R_{e}$ follows from using the bypass in the definition of effective radius, the bound by $\log^{2}n$ is straightforward from the definition of the truncated passage time. We also show in Section 2 that the effective radii have two important properties.
@@
    First, they are weakly dependent, in the sense that the event $\{R_{e}\leq t\}$ depends only on the status of edges at a distance at most $C_{*}t$ from $e$.
@@
$\sum_{e\in\gamma}\hat{R}_{e}^{2}=\sum_{M=1}^{\log^{2}n}M^{2}\sum_{e\in\gamma}{\mathbb{I}}(\hat{R}_{e}=M),$ (1.13)
and exploit the greedy lattice animals theory (see [DHS14, CN19]) to the Bernoulli random variables $({\mathbb{I}}(\hat{R}_{e}=M))_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}$ with the aid of two properties (a) and (b). In fact, we obtain a much better estimate than (1.11):
${\mathbb{P}}\Bigg{(}\sum_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}(\nabla_{e}{\mathrm{T}}_{n})^{2}\geq Cn\Bigg{)}\leq{\mathbb{P}}\Bigg{(}\sum_{e\in\gamma}\hat{R}_{e}^{2}\geq Cn\Bigg{)}\leq\exp(-n/(\log n)^{C}).$ (1.14)
@@
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}_{n}|\geq L\right)\leq c^{-1}\exp\left(-c\tfrac{L}{\log L+\log^{2}n}\right).$ (1.15)
@@
${\mathbb{P}}\Bigg{(}\sum_{e\in\Gamma}R_{e}\geq L\Bigg{)}\leq c^{-1}\exp(-cL/(\log L+\log^{2}n)).$ (1.16)
As mentioned above, achieving a large deviation of the total cost $\sum_{e\in S}R_{e}$ with $S$ a general set, as described in (1.16), is not feasible. In fact, the separable property (i) somehow strengthens the local dependent of $(R_{e})_{e\in\Gamma}$, and enables us to prove this large deviation estimate. We refer to Section 5 for the detailed proof.
Finally, let us highlight the novelty of our method. While the idea of constructing detours that bypass a given edge with reasonable length is natural and frequently applied, previous approaches often relied on multi-step renormalization processes. These processes involved constructing bypasses through a sequence of good boxes at various scales, utilizing a random number of boxes with different but deterministic radii. This sophisticated process is sensitive to the choice of scaling, see in particular [CD22, Proposition 4.5 & Remark 4.6]. In contrast, our method represents a simpler, one-step renormalization that involves a single box with a random radius. This approach allows us to achieve a nearly optimal large deviation estimate (1.14). Intuitively, we encapsulate the multi-scale analysis within the decomposition (1.13) and employ the greedy lattice animals theory to analyze each scale. This approach appears to be robust and can be applied to other problems related to controlling the effect of resampling edges in percolation, such as the Lipschitz-continuity of the time constant, as discussed in [CNN24].

### v2

@@ Abstract.
Considering supercritical Bernoulli percolation on ${\mathbb{Z}}^{d}$, Garet and Marchand [25] proved a diffusive concentration for the graph distance. In this paper, we sharpen this result by establishing the subdiffusive concentration inequality, which revisits the sublinear bound of the variance proved by Dembin [18] as a consequence. Our approach is inspired by similar work in First-passage percolation [4, 16], combined with new tools to address the challenge posed by the infinite weight of the model. These tools, including the notion of effective radius and its properties, enable a simple one-step renormalization process as a systematic method of managing the effects of resampling edges.
V.H. Can: Institute of Mathematics, Vietnam Academy of Science and Technology, 18 Hoang Quoc Viet, Cau Giay, Hanoi, Vietnam
V.Q. Nguyen: Institute of Mathematics, Vietnam Academy of Science and Technology, 18 Hoang Quoc Viet, Cau Giay, Hanoi, Vietnam
@@ 1.1. Model and main result
Bernoulli percolation is a simple but well-known probabilistic model for porous material introduced by Broadbent and Hammersley [7]. Let $d\geq 2$ and ${\mathcal{E}}({\mathbb{Z}}^{d})$} be the set of the edges $e=(x,y)$ of endpoints $x=(x_{1},\ldots,x_{d}),y=(y_{1},\ldots,y_{d})\in{\mathbb{Z}}^{d}$ such that $\|x-y\|_{1}:=\sum_{i=1}^{d}|x_{i}-y_{i}|=1$. Given the parameter $p\in(0,1)$, we let each edge $e\in{{\mathcal{E}}({\mathbb{Z}}^{d})}$ be open with probability $p$ and closed otherwise, independently of the state of other edges. The phase transition of the model has been well-known since 1960s. In particular, it has been shown that there exists a critical parameter $p_{c}(d)\in(0,1)$ such that there is almost surely a unique infinite open cluster $\mathcal{C}$$_{\infty}$ if $p>p_{c}(d)$, whereas all open clusters are finite if $p<p_{c}(d)$. We refer to the book of Grimmett [27] and the reviews of Duminil-Copin [21] for classical results and recent developments in percolation theory. In this paper, we are interested in the chemical distance that is the graph distance of infinite cluster in the supercritical regime.
@@
[…]
When $U={\mathbb{Z}}^{d}$, we simply write $\mathrm{D}$} for ${\mathrm{D}}^{{\mathbb{Z}}^{d}}$ and we write ${\mathrm{D}}(x,y)$ for ${\mathrm{D}}(\{x\},\{y\})$. Since this graph distance between two disconnected vertices is not well defined, we consider a regularized version as follows. Let $x\in{\mathbb{Z}}^{d}$, we denote by $x^{*}$ the closest point to $x$ in $\mathcal{C}$$_{\infty}$ (in $\|.\|_{\infty}$ distance) with a deterministic rule breaking ties, and call it the regularized point of $x$. The graph distance is then defined as: for $x,y\in{\mathbb{Z}}^{d}$,
[…]
@@
The linear growth of ${\mathrm{D}}^{*}_{n}$ was described by Garet and Marchand [23]: for any $p>p_{c}(d)$, there exists a constant $\mu_{p}({\mathbf{e}}_{1})\in(0,\infty)$ such that,
$\lim_{n\to\infty}\dfrac{{\mathrm{D}}^{*}_{n}}{n}=\mu_{p}({\mathbf{e}}_{1})\quad\text{a.e and in }L^{1}.$ (1)
The value $\mu_{p}({\mathbf{e}}_{1})$ is called the time constant. The regularity of the time constant (Lipschitz-continuity of the time constant with respect to $p$) was proved in [10] and revisited in [9]. Naturally, the next question we are interested in is the fluctuation and deviation of the graph distance. The large deviation of ${\mathrm{D}}^{*}_{n}$ has been well studied since 1990s, see, for example, [3, 24, 20]. The moderate deviation of ${\mathrm{D}}^{*}_{n}$ (or precisely the concentration with diffusive scale) was established by Garet and Marchand (see [25, Theorem 1.2]), using concentration inequalities of Boucheron, Lugosi, and Massart: for each $C>0$, there exists some constant $c>0$ such that for all $\lambda\in[C\log n,\sqrt{n}$],
$\mathbb{P}[|{\mathrm{D}}^{*}_{n}-{\mathbb{E}}[{\mathrm{D}}^{*}_{n}]|\geq\sqrt{n}\lambda]\leq c^{-1}\exp(-c\lambda).$ (2)
Recently, Dembin gave a sublinear bound on the variance of ${\mathrm{D}}^{*}_{n}$ (see [18, Theorem 1.1]): there exists a positive constant $C$ such that
$\operatorname{Var}[{\mathrm{D}}^{*}_{n}]\leq C\frac{n}{\log n}.$ (3)
The main result of our paper is to sharpen the moderate deviation (2) by establishing a sub-diffusive concentration of ${\mathrm{D}}^{*}_{n}$.
@@
Let $p>p_{c}(d)$. There exists a positive constant $c>0$ such that for all $n\in{\mathbb{N}}_{\geq 2}$ and $\kappa\geq 0$,
$\mathbb{P}\left(|{\mathrm{D}}^{*}_{n}-\mathbb{E}[{\mathrm{D}}^{*}_{n}]|\geq\sqrt{\frac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa).$ (4)
Consequently, the sublinear bound of the variance (3) holds.
@@
If $n$ is small and $\kappa<1$, we can choose $c$ sufficiently small so that (4) holds trivially. Throughout the rest of this paper, we therefore focus on the case where $n$ is sufficiently large and $\kappa\geq 1$.
@@
${\mathrm{T}}^{U}(x,y):=\inf_{\begin{subarray}{c}\gamma:x\to y\\ \gamma\subset U\end{subarray}}\sum_{e\in\gamma}t_{e},$ (5)
@@
${\mathrm{T}}^{*}(x,y):={\mathrm{T}}(x^{*},y^{*}),$ (6)
where $z^{*}$ denotes the closest point to $z$ in the infinite cluster of finite weight edges with a deterministic rule to break ties. Then the supercritical Bernoulli percolation can be referred as a particular case of FPP with distribution $\zeta=\zeta_{p}=p\delta_{1}+(1-p)\delta_{\infty}$.
The convergence of the scaled passage time in probability to time constant was obtained by Cerf and Théret [11, Theorem 4], without any moment assumption: there exists a constant $\mu_{\zeta}(\mathbf{e}_{1})\in[0,\infty)$ such that
$\lim_{n\to\infty}\frac{{\mathrm{T}}^{*}(0,n\mathbf{e}_{1})}{n}=\mu_{\zeta}(\mathbf{e}_{1})\qquad\text{ in probability}.$ (7)
In [23, Remark 1], Garet and Marchand proved that if ${\mathbb{E}}[t_{e}^{2+\varepsilon}{\mathbb{I}}(t_{e}<\infty)]<\infty$ with some $\varepsilon>0$, then the convergence in (7) holds true almost surely and in $L^{1}$. The continuity property of the time constant with respect to $\zeta$ was first shown by Cox and Kesten in [14, 15, 28] for FPP with finite weights. This was later extended for FPP with possibly infinite weights by Garet, Marchand, Procaccia, and Théret in [26]. It then has been proved by Damron, Hanson, and Sosoe in [17] that if $\zeta([0,\infty))=1$ and ${\mathbb{E}}[t_{e}^{2}\log_{+}t_{e}]<\infty$, then the sublinear variance holds:
$\operatorname{Var}[{\mathrm{T}}(0,n{\mathbf{e}}_{1})]\leq C\frac{n}{\log n}.$ (8)
The sublinearity of variance is a particular case of superconcentration, a phenomenon coined by Chatterjee [12] describing the situation that usual concentration tools such as Poincaré or Erfon-Stein inequality give suboptimal bounds. In fact, in some particular models as Gaussian disordered systems [12], he also established a connection between superconcentration and a chaotic phenomenon of the ground states. However, understanding these notions beyond the Gaussian realm remains quite limited. Recently, Ahlberg, Deijfen, and Sfragara [1] showed that this deep relation holds true in the context of First-passage percolation when the weight distribution $\zeta$ has a finite second moment. Other examples of the superconcentration phenomena can be found in [6, 13, 19].
@@ 1.2. Method of the proof
The general strategy of proving the subdiffusive concentration for FPP has been shown in [4] and [16]. First, by general bounds, the subdiffusive concentration of ${\mathrm{T}}(0,n{\mathbf{e}}_{1})$ can be reduced to some estimates of the variance of exponential functionals of ${\mathrm{T}}(0,n{\mathbf{e}}_{1})$. Next, to bound the variance of these exponential terms, one could use the entropy inequalities combining with the geometric averaging trick of Benjamini, Kalai, Schramm introduced in their highly influential paper [5]. Moreover, in FPP [16], the key for variance or entropy bounds is to effectively control for the effect of resampling an edge weight, and the moment conditions play a crucial role in these controls. However, in the context of the graph distance in Bernoulli percolation, closing an edge on the geodesic can have a significant impact on the graph distance due to the possibility of infinite edge-weight values. By introducing the notion effective radius and exploiting the greedy lattice animals, we provide a simple one-step renormalization process as a systematic way to manage the effect of resampling edges. Let us now explain our proof in more detail.
@@
There exists a positive constant $c$ such that for all $n\in{\mathbb{N}}_{\geq 2}$ and $\kappa\geq 0$,
$\mathbb{P}\left(|{{\mathrm{T}}_{n}}-\mathbb{E}[{{\mathrm{T}}_{n}}]|\geq\sqrt{\tfrac{n}{\log n}}\kappa\right)\leq c^{-1}\exp(-c\kappa).$ (9)
Our argument initially follows the common scheme as in [4, 16]. Considering ${\mathrm{F}}_{m}$ a spatial average of ${\mathrm{T}}_{n}$ defined in (44) (inspired by Benjamini, Kalai, and Schramm in [5]), the proof of (9) can be transfered to the following. There exists a constant $c>0$ such that with $K=\dfrac{cn}{\log n}$,
$\forall\,|\lambda|<\dfrac{1}{\sqrt{K}},\quad\operatorname{Var}\left[e^{\lambda{{\mathrm{F}}_{m}}}\right]\leq K\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]<\infty.$ (10)
Using Falik-Samorodnitsky inequality (Lemma 4.1), the problem (10) is then reduced to understand well the behavior of $G=e^{\lambda{\mathrm{F}}_{m}}$ when flipping the state of edges in ${\mathcal{E}}({\mathbb{Z}}^{d})$}. In fact, we shall show in Section 4 that the above variance bound is closely related to a large deviation estimate
${\mathbb{P}}\Bigg{(}\sum_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}(\nabla_{e}{\mathrm{T}}_{n})^{2}\geq Cn\Bigg{)}\leq\exp(-\sqrt{n}(\log n)^{C}),\qquad\nabla_{e}{\mathrm{T}}_{n}={\mathrm{T}}_{n}(\log^{2}n,t_{e^{c}})-{\mathrm{T}}_{n}(1,t_{e^{c}}),$ (11)
@@
$|\nabla_{e}{\mathrm{T}}_{n}|\leq\hat{R}_{e}{\mathbb{I}}(e\in\gamma),\quad\hat{R}_{e}=\min\{C_{*}R_{e},\log^{2}n\},$ (12)
with $C_{*}$ being a constant introduced in (18) and $\gamma$ is a geodesic of ${\mathrm{T}}_{n}$. While the bound by $C_{*}R_{e}$ follows from using the bypass in the definition of effective radius, the bound by $\log^{2}n$ is straightforward from the definition of the truncated passage time. We also show in Section 2 that the effective radii have two important properties.
@@
    First, they are weakly dependent in the sense that the event $\{R_{e}\leq t\}$ depends only on the status of edges at a distance at most $C_{*}t$ from $e$.
@@
$\sum_{e\in\gamma}\hat{R}_{e}^{2}=\sum_{M=1}^{\lfloor\log^{2}n\rfloor}M^{2}\sum_{e\in\gamma}{\mathbb{I}}(\hat{R}_{e}=M),$ (13)
and exploit the greedy lattice animals theory (see [16, 8]) to the Bernoulli random variables $({\mathbb{I}}(\hat{R}_{e}=M))_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}$ with the aid of two properties (a) and (b). In fact, we obtain a much better estimate than (11):
${\mathbb{P}}\Bigg{(}\sum_{e\in{\mathcal{E}}({\mathbb{Z}}^{d})}(\nabla_{e}{\mathrm{T}}_{n})^{2}\geq Cn\Bigg{)}\leq{\mathbb{P}}\Bigg{(}\sum_{e\in\gamma}\hat{R}_{e}^{2}\geq Cn\Bigg{)}\leq\exp(-n/(\log n)^{C}).$ (14)
@@
$\mathbb{P}\left(|{{\mathrm{D}}^{*}_{n}}-{\mathrm{T}}_{n}|\geq L\right)\leq c^{-1}\exp\left(-c\tfrac{L}{\log L+\log^{2}n}\right).$ (15)
@@
${\mathbb{P}}\Bigg{(}\sum_{e\in\Gamma}R_{e}\geq L\Bigg{)}\leq c^{-1}\exp\left(-c\tfrac{L}{\log L+\log^{2}n}\right).$ (16)
As mentioned above, achieving a large deviation of the total cost $\sum_{e\in S}R_{e}$ with $S$ a general set, as described in (16), is not feasible. In fact, the separable property (i) somehow strengthens the local dependent of $(R_{e})_{e\in\Gamma}$, and enables us to prove this large deviation estimate. We refer to Section 5 for the detailed proof.
We conclude by emphasizing the novel aspects of our approach. While the construction of detours that bypass a given edge with controlled length is a natural and commonly employed idea, existing methods often rely on intricate multi-step renormalization schemes. These typically involve constructing bypasses through sequences of good boxes across multiple scales, using a random number of boxes with different but deterministic radii. In contrast, our method introduces a simplified one-step renormalization procedure, utilizing a single box with a random radius. This streamlined construction enables us to derive a nearly optimal large deviation estimate (14). Intuitively, we encapsulate the multi-scale analysis within the decomposition (13) and leverage the greedy lattice animals theory to analyze each scale. Our method exhibits robustness and is potentially applicable to a broader class of problems involving the control of resampling effects in percolation, such as the Lipschitz continuity of the time constant, as discussed in [9].